# Cislunar Constellation Design for Space Situational Awareness with Time-Expanded Facility Location Problem

### v1

@@ Abstract
This work proposes a facility location problem framework for designing satellite constellations to provide Cislunar Space Situational Awareness (CSSA). The problem seeks to find the optimal locations of space-based observers to monitor a user-defined region of cislunar space, represented by a finite set of targets with potentially time-varying observation requirements. The optimization problem is posed as a time-expanded facility location problem (FLP); unlike a traditional FLP that focuses on assigning facilities to clients, the proposed formulation focuses on assigning observers to specific pointing directions from a discretized set. To aid with the solving process of the resulting large-scale binary linear program, a Lagrangean method (LM) based on constraint relaxation and custom heuristics is provided. The performance of the proposed formulation and method is demonstrated with several case studies that obtained CSSA constellations for monitoring the lunar sphere of influence, the cislunar Cone of Shame, and a transit window for low-energy transfers located in the Earth-Moon L2 neck region. The proposed problem formulation, along with the LM, is demonstrated to enable a fast assessment of near-optimal CSSA constellations.
@@
$a_{\rm diff}$  = target diffusion reflectance
$a_{\rm spec}$  = target specular reflectance
@@ 1 Introduction
Cislunar space is today’s most active frontier in space exploration. In the first two months of 2024 alone, there have been two successful lunar landings, namely JAXA’s SLIM and Intuitive Machines’ Odysseus missions, and multiple more entities have planned launch before the end of the year. As activity in the lunar vicinity from both government and commercial players sees an unprecedented increase, the safe operation of cislunar assets necessitates a reliable infrastructure for monitoring activities of both cooperative and non-cooperative targets. Driven by this need, interest in cislunar space situational awareness (CSSA) has also seen a significant increase [1, 2, 3, 4, 5, 6]. Several studies have considered both monitoring vast cislunar regions [7, 8, 9, 10, 11, 12, 2, 13], as well as tracking specific translunar and cislunar trajectories of interest [14, 15, 16, 17]. In particular, Wilmer et al [18, 19] and Dahlke et al [20] studied the use of libration point orbits (LPOs) for monitoring targets also located in L1 and L2 LPOs; notably, Wilmer and Bettinger [18] looked at monitoring the near-rectilinear halo orbit (NRHO), which is the planned hosting orbit for the Lunar Gateway. Finally, Patel et al [17] looked at monitoring transits from an L1 LPO to GEO.
The capability to rapidly explore the CSSA constellation design space is critical as it allows decision-makers to assess the extent to which CSSA may be conducted for a given set of monitoring demands and a given set of specifications for the observing spacecraft system. Several factors complicate the CSSA problem; one major contributor is the large number of design parameters, such as the number and location of observers, specifications of their optical sensors, requirements for the observation, and duration of the SSA activity. Another factor is the highly nonlinear dynamics and time-dependent observation conditions, which renders traditional orbital elements-based constellation design approaches inapplicable. These factors make the design of CSSA constellations complicated; to date, several works have conducted simulation-based studies on specific CSSA constellations [8, 21, 20, 5]. Vendl et al [2] presented a parametric comparison of candidate LPOs for CSSA, providing insights into the observation conditions offered by each orbit. Building on these efforts, Visonneau et al [13] posed the design as a dual-objective minimization problem, which was tackled by a variable-length genetic algorithm (GA). This formulation wrapped a simulation as part of its objective function, thereby allowing it to handle any arbitrary level of complexity in the model. Clareson et al [22] also employed a GA, and extended the realism of the simulation by considering an extended Kalman filter used to estimate the state of the observed targets. The primary drawback of these simulation-based works is the inherent, non-negligible computational cost; for example, solving a single instance of the problem in [13] took on the order of hours. Since the CSSA problem has multiple parameters and uncertainties surrounding the requirements, the slow computational performance hinders rapid assessment of the problem. Moving away from these simulation-based approaches, the work by Patel et al [17] applies the Access-profile-Pattern-Coverage (APC) decomposition scheme, originally conceived for designing Earth observation constellations [23], resulting in a binary linear program (BLP). The use of the APC allows for a much faster exploration of the trade-space, due to the existence of powerful techniques to solve BLP, primarily based on branch-and-bound (B&B) algorithms. Like many other constellation design methods, the APC decomposition scheme optimizes only the positions of the satellites; since it assumes that the satellites can view any targets within the range (i.e., no field-of-view constraints implemented), it lacks the capability to optimize the sensor pointing directions of each satellite over time considering a realistic field of view.
In broader, operations research contexts, the general class of mathematical framework for deciding the location of infrastructure assets and their assignments to clients is known as the facility location problem (FLP). FLP is a popular approach for formulating and solving location and allocation problems in terrestrial logistics applications [24, 25, 26, 27, 28, 29, 30, 31] as well as space-based applications [32, 33, 34]. In this work, we leverage the $p$-median variant of the FLP, which results in a BLP, and extend it to include the time dimension, thus accommodating the dynamic nature of motion in cislunar space. The inclusion of the time dimension into FLP is traditionally coined multi-period FLP (MPFLP) [35, 36, 37], where each “period” corresponds to the so-called planning horizon about which allocation decisions are to be made. In the context of space-based applications, the term “period” is unfortunately overloaded, as it could also denote the period of an orbit or the synodic period of the Earth-Moon-Sun system; thus, to avoid confusion with these definitions, we refer to the formulation with the time-dimension as the time-expanded $p$-median problem (TE-$p$-MP). This is also motivated by the fact that even though the MPFLP and the TE-$p$-MP have a mathematical resemblance, the physical meaning of a decision time step is drastically different: following the definition of strategic, tactical, and operational level decisions from Hax and Candea [38], the time-step in the TE-$p$-MP is at the operational level (on the order of hours), while the time-step in the MPFLP is at the strategic or tactical level (on the order of months). The MPFLP has been developed in large parts for supply chain and inventory management applications, where demands of commodities typically have predictable, seasonal fluctuations [37]. Meanwhile, to this date, the use of MPFLP in space-based applications is limited; the few exceptions include the authors’ previous works on the SSA problem [33] and on cislunar communication relays [34], where facilities have been located to discretized slots along a given set of LPOs, and allocated to observe or provide service to specific targets/clients.
In this work, we develop a TE-$p$-MP formulation for SSA. The novelty in the proposed formulation is twofold: firstly, at each time step, the allocation of facilities (i.e. observers) to clients (i.e. observation targets) is replaced with the allocation of facilities (i.e., observers) to directions to which the sensor is pointed. The idea of allocating observers to pointing directions rather than individual target(s) to be observed is akin to also solving the sensor tasking problem [10, 11, 39] and enables formulating the SSA problem with realistic constraints based on the field of view (FOV) of the observer’s optical sensor, without introducing a conservative capacity assumption on the number of targets it may observe at a given instance in time. Secondly, an objective function that accounts for the aggregated coverage of the target volume is defined; this avoids the optimization problem from favoring double-counting targets that are easier to observe at the cost of omitting ones that are harder to observe.
To solve the TE-$p$-MP, we propose a Lagrangean method (LM) based on Lagrangean relaxation along with specifically derived heuristics developed to rapidly construct feasible solutions. The combination of the TE-$p$-MP and the LM enables rapid optimization of CSSA constellations with no restriction on observer locations, generating a sensor pointing schedule for each observer. These TE-$p$-MP instances are demonstrated for SSA architectures tasked with monitoring either a static or dynamic set of predefined targets.
The remainder of this paper is organized as follows: in Section 2, we provide an overview of the dynamical system in which the problem is defined, along with definitions of LPOs and LETs. This is followed by Section 3, where the static and dynamic SSA demands are defined, along with the observation model adopted in this work. Sections 4 and 5 are used to respectively introduce the various TE-$p$-MP instances and their Lagrangean relaxation schemes. Numerical experiments using the proposed approach are shown in Section 6. Finally, Section 7 provides conclusions to this work.
2 Dynamical Systems Preliminaries
2.1 Equations of Motion
@@
In this work, the circular restricted three-body problem (CR3BP) is used to model the motion of the observer spacecraft. The frame is defined with its first axis aligned with the vector pointing from the Earth to the Moon, the third axis aligned with the angular momentum vector of the rotating motion of the two primary bodies, and the second axis completing the triad. The motion of the two primaries are assumed to be circular about their common barycenter. Let $\boldsymbol{x}\in\mathbb{R}^{6}$ denote the state of the spacecraft, composed of $\boldsymbol{x}=[\boldsymbol{r}^{T},\boldsymbol{v}^{T}]^{T}$, where $\boldsymbol{r}\in\mathbb{R}^{3}$ denote the position of the spacecraft, $\boldsymbol{v}\triangleq\dot{\boldsymbol{r}}\in\mathbb{R}^{3}$ denote the velocity of the spacecraft in the rotating frame. The CR3BP equations of motion $\boldsymbol{f}(\boldsymbol{x})$ are given by
@@
$\begin{bmatrix}\dot{\boldsymbol{r}}\\ \ddot{\boldsymbol{r}}\end{bmatrix}=\boldsymbol{f}(\boldsymbol{x})=\begin{bmatrix}\boldsymbol{v}\\ -\dfrac{\mu_{1}}{\|\boldsymbol{r}_{1}\|_{2}^{3}}\boldsymbol{r}_{1}-\dfrac{\mu_{2}}{\|\boldsymbol{r}_{2}\|_{2}^{3}}\boldsymbol{r}_{2}-\boldsymbol{\omega}\times\left(\boldsymbol{\omega}\times\boldsymbol{r}\right)-2\boldsymbol{\omega}\times\boldsymbol{v}\end{bmatrix}$ (1)
@@
where $\boldsymbol{r}_{1}\in\mathbb{R}^{3}$ and $\boldsymbol{r}_{2}\in\mathbb{R}^{3}$ are position vectors to the first and second primary bodies with gravitational parameters $\mu_{1}$ and $\mu_{2}$, and $\boldsymbol{\omega}=[0,0,\omega]^{T}$ is the angular momentum vector of the Earth-Moon rotating frame. The corresponding state-transition matrix (STM) $\Phi(t,t_{0})\in\mathbb{R}^{6\times 6}$, mapping an initial linear perturbation $\delta\boldsymbol{x}\in\mathbb{R}^{6}$ at time $t_{0}$ to time $t$, is propagated by the initial value problem
@@
We choose canonical scales such that $\mu_{1}+\mu_{2}=1$ and $\omega=1$, the system parameters are given in Table 2.
@@
2.2 Libration Point Orbit

### v2

# Cislunar Constellation Design for Space Situational Awareness with Time-Expanded p-Median Problem

@@ Abstract
Driven by the surmounting interest for a dedicated infrastructure in cislunar space, this work considers the satellite constellation design for cislunar space situational awareness (CSSA). We propose a linear programming (LP)-based formulation that simultaneously tackles the constellation design and sensor-tasking subproblems surrounding CSSA. Our approach generates constellation designs that provide coverage with considerations for the field-of-view of observers. We propose a time-expanded $p$-Median problem (TE-$p$-MP) which considers the optimal placement of $p$ space-based observers into discretized locations based on orbital slots along libration point orbits, simultaneously with observer pointing directions across discretized time. We further develop a Lagrangian method for the TE-$p$-MP, where a relaxed problem with an analytical solution is derived, and customized heuristics leveraging the orbital structure of candidate observer locations are devised. The performance of the proposed formulation is demonstrated with several case studies for CSSA constellations monitoring the cislunar Cone of Shame and a periodic time-varying transit window for low-energy transfers located in the Earth-Moon L2 neck region. The proposed problem formulation, along with the Lagrangian method, is demonstrated to enable a fast assessment of near-optimal CSSA constellations, equipping decision-makers with a critical technique for exploring the design trade space.
@@
$C_{\rm diff}$  = Lambertian diffuse reflection coefficient
@@ 1 Introduction
Cislunar exploration is one of the defining themes in space exploration of this decade. Spearheaded by programs such as Artemis and the Commercial Lunar Payload Services (CLIPS), there is a growing number of public and private sector missions to both the lunar surface and various lunar orbits. The growing traffic in cislunar space leads to an increasing need for a dedicated cislunar space situational awareness (CSSA) infrastructure - one that can detect and/or monitor the motion of both cooperative and uncooperative assets [1, 2, 3, 4, 5, 6]. In contrast to Earth orbits, CSSA is complicated by two defining characteristics: the vastness of the volume of interest to be monitored, and the higher nonlinearity of objects in cislunar vicinities. Past works have either considered monitoring a general volume of interest [7, 8, 9, 10, 11, 12, 2, 13], or focused on monitoring specific translunar and cislunar trajectories of interest [14, 15, 16, 17, 18, 19, 20]. The first characteristic renders certain parts of the volume poorly observable from Earth-based observers, thus requiring observations to be conducted from dedicated satellite constellations. The second characteristic necessitates more frequent observation and/or information of higher quality to be provided to prediction algorithms to compensate for the higher nonlinearity. These defining features of CSSA drive the need for a dedicated constellation infrastructure. For a comprehensive review of the CSSA problem, see the review paper by Baker-McEvilly et al. [6] and references therein.
At its core, the CSSA problem comprises two major problems: constellation design [1, 2, 3, 4, 5, 6, 7, 8, 9, 12, 13, 14, 15, 16, 17, 18, 19, 20] and sensor-tasking [10, 11, 21]. Constellation design consists of finding the optimally distributed network for performing SSA, while sensor-tasking consists of solving for an optimal schedule for each segment within the network to maximize some monitoring performance metric. Due to their complexity, each of these problems is commonly tackled individually: constellation design typically assumes volumes of interest can be simultaneously observed at any time as long as it is visible from an observer, while sensor-tasking assumes an already “fixed” constellation, and solves for the optimal observation schedule.
Future activities in cislunar space remain difficult to predict with certainty, and so is the exact demand for CSSA. Observer parameters, such as sensor specifications or the number of assets to be deployed, also remain uncertain. To explore the vast trade space for CSSA infrastructure across different combinations of demand and observer parameters, a mathematical framework to efficiently assess the performance of a given constellation is necessitated.
1.1 Overview of Existing Approaches for Cislunar Space Situational Awareness Problems
Previous works tackling either the constellation design or sensor-tasking problems can be categorized into either nonlinear programming (NLP)-based or linear programming (LP)-based approaches.
In an NLP-based approach, a simulation environment is used for parametric studies [7, 2, 22, 17, 5] or as part of an objective function for algorithms such as evolutionary algorithms (EA) [13, 23], Monte Carlo tree search (MCTS) [11, 24, 25], or machine learning (ML) [26, 27, 28, 29], to name a few. Due to EA and MCTS handling the objective function as a black box, any arbitrary level of fidelity may be incorporated into the simulation environment without altering the algorithm used. Meanwhile, the typically high-fidelity nature of such simulation environments results in objective functions that are computationally expensive, thereby prohibiting extensive parameter trade-off studies. Furthermore, the curse of dimensionality renders NLP-based approaches ill-suited for problem formulations involving large numbers of variables, thus necessitating alternative formulations that avoid extensive combinatorial decisions to be made.
In an LP-based approach, an LP model, commonly involving binary or integer variables, is constructed and solved [30, 21, 31, 20]. Nonlinearities surrounding either the constellation design or sensor-tasking, are precomputed to form objectives and/or constraints to the model. Mixed-integer LP (MILP) and binary LP (BLP) are well-suited for encoding combinatorial decisions, and a careful formulation can lead to problem formulations that can be efficiently solved by specialized algorithms such as branch and bound (BnB) [32, 33, 34, 35], Lagrangian method (LM) [36, 37, 38], or column generation [39, 40, 41]. Additionally, LP algorithms provide an optimality gap, which corresponds to the worst-case degradation of a prematurely terminated suboptimal solution from the predicted optimal solution. Such a gap is not readily available in NLP-based approaches such as EA, MCTS, or ML. Notable challenges surrounding LP-based approaches include the need for a reasonable discretization scheme to approximate the continuous decision space in observer placement and sensor tasking, and devising an appropriate formulation that is tractable.
1.2 Problem Definition and Proposed Approach
The constellation design and sensor-tasking subproblems of CSSA are fundamentally coupled. The sensor-tasking schedule depends on the architecture, while the actual performance of an architecture is the result of the sensor-tasking schedule. To go beyond the approximation of simultaneously monitoring an arbitrary region of space and instead consider observations restricted by the sensor’s field-of-view (FOV), one must combine the sensor-tasking of each observer into the constellation design problem.
In this work, we propose a logistics-inspired formulation to simultaneously solve the constellation design and sensor-tasking problems. Our formulation takes inspiration from the $p$-median problem, a classical formulation for studying the placement problem of facilities in terrestrial logistics. We consider the placement of observer spacecraft into orbital “slots” along libration point orbits (LPO) and their allocation to a pointing direction along which the observers monitor. To accommodate for the time-varying position of observers along their respective orbits, we augment the allocations with a time index, such that the observers are prescribed a specific pointing direction for each time step. The resulting formulation is referred to as the time-extended $p$-median problem (TE-$p$-MP). To the best of the authors’ knowledge, time-expanded $p$-median formulations as applied to space problems are limited to previous iterations of this work [30, 42]. While this work concerns CSSA, an appropriately modified TE-$p$-MP formulation can be employed for various other applications, such as communication relay, planetary defense, or in-orbit servicing, to name a few.
While generic BnB algorithms are known to adequately tackle classical $p$-median problems of moderate size, we also propose a custom LM to compute an approximate solution rapidly. LM consists of iteratively solving for an idealized optimal solution to a relaxed problem and a heuristics-based feasible solution, each time updating Lagrange multipliers used to construct the relaxed problem. In this paper, an LM with a purely analytical solution to the relaxed problem is derived by decoupling the observer-wise sensor-tasking decision. Custom heuristics for the feasible solution are designed to leverage the orbital structure of candidate observer locations. We note that the large number of variables, in the order of $10^{6}$, of the TE-$p$-MP prohibits the use of EAs. An NLP-based approach for tackling both constellation design and sensor-tasking problems could consist, for example, of formulating a black-box constellation design problem in the outer loop, with sensor-tasking handled in the inner loop by a pre-trained neural network (NN), or a policy chosen from a predefined set by the outer loop. We however opt for the proposed LP-based approach in favor of its ease in constructing a tractable formulation that encodes decision surrounding the observation schedule. With an NLP-based approach, a policy-based approach would result in a limited performance due to the inherently enforced structure by the pre-defined set of policies, while parameterizing the policy to be optimized by an outer-loop is non-trivial. Recently, Patel et al. [20] also proposed a framework for concurrent CSSA constellation design and sensor-tasking. Their approach is a hybrid between NLP and LP, where an outer-layer NLP optimizes the phasing of satellites along pre-defined orbits, and an inner-layer LP that assigns discrete targets to observers. This work differs from [20] in the following ways: our formulation optimizes not only for the phasing but also the orbits of observers from a discrete set of options; we also consider target observation by considering a sensor FOV, rather than assuming a one-to-one assignment between observers and targets; finally, our formulation is a single-layer MILP that does not rely on a black-box NLP solver. Through the proposed TE-$p$-MP together with the LM, our approach can produce a feasible candidate CSSA constellation design together with a pointing direction schedule for each observer that provides adequate coverage to a given set of demand and observer parameters.
The remainder of this paper is organized as follows: in Section 2, we provide background information on the facility location problem, the $p$-Median problem, the Lagrangian method, and the dynamical system within which the TE-$p$-MP is considered. This is followed by Section 3, where the static and dynamic SSA demands are defined, along with the observation model adopted in this work. Respectively, Sections 4 and 5 are used to introduce the various TE-$p$-MP instances and their Lagrangian relaxation schemes. Numerical experiments using the proposed approach are shown in Section 6. Finally, Section 7 provides conclusions to this work.
2 Background
We first provide background on the facility location problem (FLP), and more specifically the $p$-Median problem. These form the basis of the TE-$p$-MP formulated in Section 4 for the CSSA application. We then briefly introduce of the Lagrangian method for a generic integer LP. Finally, we provide a review of the circular restricted three-body dynamics model adopted to model the motion of observers, along with synodic resonances that exist for LPOs.
2.1 Facility Location Problem
In Operations Research, the general class of mathematical framework for deciding the location of infrastructure assets and their assignments to clients is known as the facility location problem (FLP). FLP is a popular approach for formulating and solving location and allocation problems in terrestrial logistics applications [43, 44, 45, 46, 47, 48, 49, 50] as well as space-based applications [51, 30, 42].
2.1.1 $p$-Median Problem
One variant of the FLP is the $p$-Median problem, where the number of facilities to be deployed is prescribed. Let $j=1,\ldots,n$ index candidate locations where a facility may be built, and $i=1,\ldots,m$ index the clients to be serviced. In the classical $p$-Median problem, let $X\in\mathbb{B}^{m\times n}$ and $Y\in\mathbb{B}^{n}$ denote
$Y_{j}$ =$=\begin{cases}1&\text{facility location $j$ is used}\\ 0&\text{otherwise}\end{cases}$ (1)
$X_{ij}$ =$=\begin{cases}1&\text{client $i$ is allocated to facility location $j$}\\ 0&\text{otherwise}\end{cases}$ (2)
Then, the $p$-Median problem is given by
$\min_{X,Y}\quad$ $\sum_{j=1}^{n}f_{j}Y_{j}+\sum_{i=1}^{m}\sum_{j=1}^{n}c_{ij}X_{ij}$ (3a)
s.t. $\sum_{j=1}^{n}X_{ij}=1\quad\forall i=1,\ldots,m$ (3b)
$\sum_{j=1}^{n}Y_{j}=p$ (3c)
$X_{ij}\leq Y_{j}\quad\forall i=1,\ldots,m,\,j=1,\ldots,n$ (3d)
$X_{ij},Y_{j}\in\{0,1\},\quad\forall i=1,\ldots,m,\,j=1,\ldots,n$ (3e)
where $f_{j}$ is the deployment cost of a candidate facility location $j$, and $c_{ij}$ denote the allocation cost of a facility at $j$ to client $i$. The objective (3a) is based on the cumulative sum of the deployment costs of the $p$ facilities and the total allocation cost; constraints (3b) ensure each client is serviced by exactly one facility; constraint (3c) ensures exactly $p$ facility locations are used; constraints (3d) ensure if at least one client is allocated to facility location $j$, then there is a facility at $j$; finally, constraints (3e) ensure $X_{ij}$ and $Y_{j}$ are binary. The TE-$p$-MP formulated in this work is based on (3), but with appropriate modifications made to (1) accommodate an objective function based on cumulative target coverage and to (2) include time indices.
2.1.2 Including Time in the Facility Location Problem
The inclusion of the time dimension into FLP is traditionally coined multi-period FLP (MPFLP) [52, 53, 54], where each “period” corresponds to the so-called planning horizon about which allocation decisions are to be made. In the context of space-based applications, the term “period” is unfortunately overloaded, as it could also denote the period of an orbit or the synodic period of the Earth-Moon-Sun system. To avoid confusion with these definitions, we refer to the proposed formulation with the time dimension as the time-expanded $p$-median problem (TE-$p$-MP). The naming convention adopted in this work is also motivated by the fact that even though the MPFLP and the TE-$p$-MP have a mathematical resemblance, the physical meaning of a decision time step is drastically different: following the definition of strategic, tactical, and operational level decisions from Hax and Candea [55], the time-step in the TE-$p$-MP is at the operational level (on the order of hours), while the time-step in the MPFLP is at the strategic or tactical level (on the order of months). The MPFLP has been developed in large parts for supply chain and inventory management applications, where demands of commodities typically have predictable, seasonal fluctuations [54]. Meanwhile, to this date, the use of MPFLP in space-based applications is limited; the few exceptions include the authors’ previous works on the SSA problem [30] and on cislunar communication relays [42], where facilities have been located to discretized slots along a given set of LPOs, and allocated to observe or provide service to specific targets/clients.
2.2 Lagrangian Method for Integer Linear Program
One popular approach to solving complex binary and integer linear programs (ILP/BLP) is the Lagrangian method (LM), which uses Lagrangian multipliers to relax the original problem and iteratively update the multipliers.
Consider a generic ILP given by
$\max_{\boldsymbol{z}}$ $\quad\boldsymbol{f}^{T}\boldsymbol{z}$ (4a)
s.t. $\boldsymbol{A}_{s}\boldsymbol{z}\geq\boldsymbol{b}_{s}$ (4b)
$\boldsymbol{A}_{c}\boldsymbol{z}\geq\boldsymbol{b}_{c}$ (4c)
$\boldsymbol{z}\in\mathbb{Z}_{\geq 0}$ (4d)
with an optimal solution $Z^{*}$, where constraints (4b) are simple, and constraints (4c) are complicating constraints. Determining a given constraint is simple or complicating is problem-specific and requires careful judgment about the problem at hand. The fundamental intuition behind LM is to remove complicating constraints by adjoining them using Lagrange multipliers into the objective, in such a way that the original ILP may be decoupled into smaller subproblems that are easier to solve. Let $\boldsymbol{\lambda}\geq\boldsymbol{0}$ denote Lagrangian multipliers for each complicating constraint. The relaxed problem is given by
$\max_{\boldsymbol{z}}$ $\quad\boldsymbol{f}^{T}\boldsymbol{z}-\boldsymbol{\lambda}^{T}\left(\boldsymbol{b}_{c}-\boldsymbol{A}_{c}\boldsymbol{z}\right)$ (5)
s.t. (4b), (4d)
whose solution is denoted by $Z_{\rm relax}$. Since problem (5) is a relaxation of (4), the solution to (5) serves as an upper-bound to the original ILP (4), which however may not be a feasible solution (4). Thus, simultaneously, a feasible lower bound solution is reconstructed from $Z_{\rm feas}$ leveraging problem-specific heuristics, such that
$Z_{\rm feas}\leq Z^{*}\leq Z_{\rm relax}$ (6)
where $Z_{\rm relax}-Z_{\rm feas}$ represents the optimality gap.
At each iteration, the LM iteratively updates $\boldsymbol{\lambda}$ to solve for
@@ 2.2 Lagrangian Method for Integer Linear Program
Since (7) is a piecewise linear concave function with respect to $\boldsymbol{\lambda}$ [48], a commonly adopted approach to solve for iterates of $\boldsymbol{\lambda}$ is the subgradient method [56]. At a given iteration $h$, $\boldsymbol{\lambda}$ is updated with
$\boldsymbol{\lambda}^{(h+1)}$ $=\boldsymbol{\lambda}^{(h)}+s^{(h)}\left(\boldsymbol{b}_{c}-\boldsymbol{A}_{c}\bar{\boldsymbol{z}}^{(h)}\right)$ (8)
where $s^{(h)}$ is the step-size given by
$s^{(h)}=\dfrac{\mu^{(h)}\left(Z_{\mathrm{relax}}^{(h)}-Z_{\mathrm{feas}}^{(h)}\right)}{\sum{\max(0,\boldsymbol{b}_{c}-\boldsymbol{A}_{c}\bar{\boldsymbol{z}}^{(h)})}}$ (9)
In equation (9), $\mu^{(h)}$ is a hyperparameter for scaling the step-size, which is reduced if no reduction of the gap between the lower-bound and upper-bound solutions is observed; initially, $\mu^{(0)}=2.0$ is typically chosen when employing the subgradient update, with a step reduction factor of 0.5 [48].
2.3 Dynamical Systems Preliminaries
2.3.1 Equations of Motion
@@
In this work, the circular restricted three-body problem (CR3BP) is used to model the motion of the observer spacecraft. In the CR3BP, the motion of a spacecraft of negligible mass is studied under the gravitational forces of the Earth and the Moon. The two celestial bodies are assumed to be co-rotating about their common barycenter along a circular orbit. Following standard notation, we consider the CR3BP dynamics represented in the Earth-Moon rotating frame, defined with its center at the barycenter, the $x$-axis along the Earth-Moon direction, and the $z$-axis aligned with the angular momentum vector of the Earth and the Moon. Let $\boldsymbol{x}\in\mathbb{R}^{6}$ denote the state of the spacecraft, composed of $\boldsymbol{x}=[\boldsymbol{r}^{T},\boldsymbol{v}^{T}]^{T}$, where $\boldsymbol{r}\in\mathbb{R}^{3}$ denote the position of the spacecraft, $\boldsymbol{v}\triangleq\dot{\boldsymbol{r}}\in\mathbb{R}^{3}$ denote the velocity of the spacecraft in the rotating frame. The CR3BP equations of motion $\boldsymbol{f}(\boldsymbol{x})$ are given by
@@
$\begin{bmatrix}\dot{\boldsymbol{r}}\\ \dot{\boldsymbol{v}}\end{bmatrix}=\boldsymbol{f}(\boldsymbol{x})=\begin{bmatrix}\boldsymbol{v}\\ -\dfrac{\mu_{1}}{\|\boldsymbol{r}_{1}\|_{2}^{3}}\boldsymbol{r}_{1}-\dfrac{\mu_{2}}{\|\boldsymbol{r}_{2}\|_{2}^{3}}\boldsymbol{r}_{2}-\boldsymbol{\omega}\times\left(\boldsymbol{\omega}\times\boldsymbol{r}\right)-2\boldsymbol{\omega}\times\boldsymbol{v}\end{bmatrix}$ (10)
where $\boldsymbol{r}_{1}\in\mathbb{R}^{3}$ and $\boldsymbol{r}_{2}\in\mathbb{R}^{3}$ are position vectors of the Earth and the Moon with gravitational parameters $\mu_{1}$ and $\mu_{2}$, and $\boldsymbol{\omega}=[0,0,\omega]^{T}$ is the angular velocity vector of the Earth-Moon rotating frame. Making use of canonical scales defined in Table 2, $\omega=1$, $\mu_{1}=1-\mu_{2}$, and $\boldsymbol{r}_{1}$ and $\boldsymbol{r}_{2}$ are given by
$\boldsymbol{r}_{1}=[x+\mu_{2},y,z]^{T},\quad\boldsymbol{r}_{2}=[x-1+\mu_{2},y,z]^{T}$ (11)
@@
The corresponding state-transition matrix (STM) $\Phi(t,t_{0})\in\mathbb{R}^{6\times 6}$, mapping an initial linear perturbation $\delta\boldsymbol{x}\in\mathbb{R}^{6}$ at time $t_{0}$ to time $t$, is propagated by the initial value problem
@@
2.3.2 Libration Point Orbit